# CogCanvas: Compression-Resistant Cognitive Artifacts for Long LLM Conversations

### v1

@@ 5  Conclusion
CogCanvas demonstrates that training-free, verbatim-grounded memory can achieve substantial improvements over standard baselines, particularly on temporal reasoning (+530% vs. GraphRAG, +239% vs. RAG) and multi-hop questions (+41.0pp vs. GraphRAG). While absolute performance remains below heavily-optimized fine-tuned approaches (EverMemOS: $\sim$92%, EMem: $\sim$78%), our work provides practitioners with an immediately-deployable solution that requires no training infrastructure.
We release our code and benchmarks to facilitate future work on compression-tolerant memory systems. [^2] https://github.com/tao-hpu/cog-canvas Future directions include adaptive threshold learning, cross-lingual artifact extraction, and hybrid strategies that combine structured artifacts with selective summarization. The whiteboard effect points toward a general principle for LLM memory systems: rather than compressing information into lossy summaries, identify and preserve the structured artifacts that anchor understanding.
@@ 6  Limitations
Design trade-offs. CogCanvas’s training-free design prioritizes immediate deployability over maximum accuracy. Our 34.7% overall accuracy on LoCoMo, while substantially outperforming standard baselines (RAG: 25.6%, GraphRAG: 13.7%), remains below heavily-optimized fine-tuned approaches (EverMemOS: 92.3%, EMem: 78%). This reflects our deliberate choice: we optimize for accessibility (any LLM backend, no training required) rather than peak performance.
@@
Single-hop trade-off. CogCanvas’s structured extraction optimizes for temporal and multi-hop reasoning while achieving competitive single-hop retrieval (35.5% vs. RAG’s 39.7%). For applications requiring primarily simple fact lookup, RAG may still have a slight edge; CogCanvas excels across all question types with particular strength in complex reasoning scenarios.
@@ 7  Future Work
Single-hop optimization. While our temporal-aware construction dramatically improves temporal reasoning (+530% vs. GraphRAG, +239% vs. RAG), achieving competitive performance across all question types, future work could explore adaptive strategies that further optimize for simple fact lookup while maintaining strong complex reasoning capabilities.
@@ Appendix C Case Study: Qualitative Analysis
Q: “When did Caroline attend the LGBTQ support group?” CogCanvas correctly retrieved the temporal context through our temporal-aware graph construction, achieving 27.1% temporal accuracy—dramatically outperforming baselines (4–6%).

### v2

@@ 5  Conclusion
Our evaluation on the LoCoMo benchmark (all 10 conversations from the ACL 2024 release) demonstrates that CogCanvas achieves the highest overall accuracy among training-free methods (32.4%), outperforming RAG (24.6%) by +7.8pp. CogCanvas leads across all question types: +20.6pp on temporal reasoning (32.7% vs. 12.1%), +1.1pp on multi-hop questions (41.7% vs. 40.6%), and +2.0pp on single-hop retrieval (26.6% vs. 24.6%). This validates our core hypothesis: structured artifact preservation with enhanced retrieval provides consistent advantages for long-context memory tasks.
While stateful systems like Letta [letta2024memory] achieve higher absolute accuracy (74.0%) through persistent memory infrastructure, our work demonstrates that training-free, stateless artifact extraction can provide a competitive alternative for scenarios prioritizing deployment simplicity, particularly excelling in temporal and multi-hop reasoning tasks.
We release our code and benchmarks to facilitate future work on compression-tolerant memory systems. [^4] https://github.com/tao-hpu/cog-canvas Future directions include adaptive threshold learning, cross-lingual artifact extraction, and hybrid strategies that combine structured artifacts with selective summarization. The whiteboard effect points toward a general principle: for complex reasoning over long conversations, preserving structured cognitive artifacts outperforms both lossy compression and simple retrieval.
@@ 6  Limitations
Design trade-offs. CogCanvas’s training-free design prioritizes immediate deployability over maximum accuracy. Our 32.4% overall accuracy on LoCoMo outperforms RAG (24.6%) by +7.8pp, with both substantially outperforming GraphRAG (10.6%) and summarization (5.6%). All training-free methods remain below heavily-optimized fine-tuned approaches (EverMemOS: 92.3%, EMem: 78%). This reflects our deliberate choice: we optimize for accessibility (any LLM backend, no training required) rather than peak performance.
GraphRAG baseline configuration. Our GraphRAG experiments use the official Microsoft implementation with default parameters (local search, chunk size=800, community_level=2). We did not tune entity extraction prompts, chunk strategies, or community detection for conversational memory. GraphRAG’s low performance (10.6%) may partially reflect suboptimal configuration rather than fundamental limitations. However, our core finding—that verbatim grounding outperforms lossy summarization—is validated by comparison with the Summarization baseline (5.6%), which shares GraphRAG’s compression-based philosophy.
@@
Enhanced retrieval parameters. CogCanvas uses increased retrieval parameters (top_k=20, graph_hops=4) to achieve consistent advantages across all question types. This may increase computational overhead compared to simpler RAG configurations, though it remains modest compared to fine-tuning pipelines.
@@ 7  Future Work
Further optimization. While CogCanvas now leads across all question types (temporal: +20.6pp, multi-hop: +1.1pp, single-hop: +2.0pp vs. RAG), future work could explore adaptive retrieval parameters that dynamically adjust based on query complexity.
@@ Appendix C Case Study: Qualitative Analysis
Q: “When did Caroline attend the LGBTQ support group?” CogCanvas correctly retrieved the temporal context through our verbatim-grounded extraction, achieving 32.7% temporal accuracy—substantially outperforming RAG (12.1%) and GraphRAG (3.1%).
@@ Appendix F Threshold Sensitivity
Appendix G Extended Related Work
This section provides detailed discussion of related work summarized in Section 2.
G.1  Context Window Management
G.2  Memory-Augmented Language Models
MemGPT [packer2023memgpt] pioneered a hierarchical memory system inspired by operating system design, introducing virtual context management where LLMs perform explicit memory management operations (moving data between “fast” and “slow” memory tiers). The original MemGPT project has since evolved into Letta, [^5] https://letta.ai a production-ready framework that extends the core architecture with filesystem-based memory storage, PostgreSQL persistence, and enhanced tool integration. Letta’s recent LoCoMo evaluation achieved 74.0% accuracy using GPT-4o mini [letta2024memory] by storing conversation histories as files with semantic search capabilities. While this demonstrates the power of stateful, database-backed memory systems, it requires persistent infrastructure (PostgreSQL, Docker) that introduces deployment complexity. This architectural approach represents a fundamentally different paradigm from our stateless, inference-time retrieval.
G.3  Graph-Based Knowledge Organization
G.4  Conversation Summarization